OpenStudy (anonymous):

Find the arc length of the curve x=1/3(y^2+2)^3/2 from y=0 to y=1

OpenStudy (turingtest):

where are you stuck?

OpenStudy (anonymous):

After I input it into the formula, I'm having trouble simplifying it

OpenStudy (turingtest):

can you show what you have so far?

OpenStudy (anonymous):

integrate (1/2) sqrt(6 + y^2) from 0 to1

OpenStudy (anonymous):

@TuringTest sqrt 1+ (y(y^2+2)^1/2)^2

OpenStudy (anonymous):

@pi3 where did the 6 come from?

OpenStudy (turingtest):

simplify \((y(y^2+2)^{1/2})^2\) by using the rule \((x^a)^b=x^{ab}\)

OpenStudy (turingtest):

alongside the rule \((xy)^a=x^ay^a\)

OpenStudy (anonymous):

okay so i use that rule with the (y^2+2)^1/2 first?

OpenStudy (turingtest):

the order doesn't matter

OpenStudy (turingtest):

\[(y(y^2+2)^{1/2})^2=(y)^2[(y^2+1)^{1/2}]^2=?\]

OpenStudy (anonymous):

where did the 1 come from?

OpenStudy (turingtest):

should have been a 2, sorry

OpenStudy (turingtest):

I just distributed the exponent

OpenStudy (turingtest):

\[(y(y^2+2)^{1/2})^2=(y)^2[(y^2+2)^{1/2}]^2=?\]

OpenStudy (anonymous):

oh okay lol

OpenStudy (anonymous):

so i would get (y)^2 [(y^2+2)]^5/2?

OpenStudy (turingtest):

not quite\[[(y+2)^a]^b=(y+2)^{ab}\]what are a and b in your case?

OpenStudy (anonymous):

they are 1/2 and 2, are they suppose to multiply?

OpenStudy (turingtest):

yes